nformation from the financial statements of Ames Fabricators, Inc., included the following:
  

December 31
2021 2020
Common shares 100,000 100,000
Convertible preferred shares (convertible into 32,000 shares of common) 12,000 12,000
8% convertible bonds (convertible into 30,000 shares of common) $ 1,000,000 $ 1,000,000


Ames’s net income for the year ended December 31, 2021, is $500,000. The income tax rate is 25%. Ames paid dividends of $5 per share on its preferred stock during 2021.

Required:

Compute basic and diluted earnings per share for the year ended December 31, 2021. (Enter your answers in thousands. Do not round intermediate calculations.)
Answer #1

Basic earning per share = (Net income-Preferred dividend)/Weighted average share outstanding

= (500000-12000*5)/100000

Basic earning per share = 4.40 per share

Diluted earning per share = Adjusted net income/adjusted share = (500000+60000)/162000 = 3.46 per share